OpenStudy (dan815):

a+b+c+d=k a,b,c,d>=1 can you find a formula for the number of integer solutions as a function of k?

OpenStudy (anonymous):

It looks like repeated choose operation.

ganeshie8 (ganeshie8):

\[\binom{k-1}{3}\]

OpenStudy (dan815):

how did u get that so fast lol

ganeshie8 (ganeshie8):

it is a good ol stars and bars problem \(k\) stars and \(3\) bars

OpenStudy (dan815):

wait wont u have a 0 in the middle with that form

OpenStudy (dan815):

oh did u subtract 3 spots

OpenStudy (dan815):

for the 3 zeros not allowed

ganeshie8 (ganeshie8):

\[\star\star|\star\star|\star|\star\star\]

OpenStudy (dan815):

like if zeros are allowed it would be K+3 choose 3 right

ganeshie8 (ganeshie8):

Yep

OpenStudy (dan815):

how come this works, so u subtract 4 spots? for the 4 zeros no allowed

ganeshie8 (ganeshie8):

Firstly, notice that those 3 bars divide the 7 stars into 4 pieces

OpenStudy (dan815):

right

ganeshie8 (ganeshie8):

You want to have at least 1 star in each piece, yes ?

OpenStudy (dan815):

right

OpenStudy (dan815):

okay i see now that limits the total number of places

OpenStudy (dan815):

ahhh

ganeshie8 (ganeshie8):

thats it! so only the places between stars are allowed

OpenStudy (dan815):

dang that was simple lol

OpenStudy (dan815):

i got to this point with some other little trick lol

OpenStudy (dan815):

like u know if u look at the difference of the solutions that has a stars and bar solution again allowing 0

ganeshie8 (ganeshie8):

solving a+b+c+d = k over nonnegative integers ?

OpenStudy (dan815):

kinds cool that exists

OpenStudy (dan815):

yeah like umm

OpenStudy (dan815):

say we were looking at a+b+c=7 then the simple case 1+1+5=7 now i looked at their differences

OpenStudy (dan815):